(Q) if you ride a bicycle 5 MPH, How many miles will you ride in 45 minutes?
(A) can you tell me if 4 1/7 is correct.

Answer by mananth(16946)

d=rt
r=5mph
t =3/4 hours
d= 5*3/4
d=15/4
=3 3/4
...
m.ananth@hotmail.ca